OpenStudy (ellamoyseyuk):

the volume of a sphere is 80 pi cm3. What is its surface area to the nearest whole number?

OpenStudy (ellamoyseyuk):

how would i solve this i have no clue. @

OpenStudy (robtobey2):

First calculate the sphere's radius by solving the following for r.

OpenStudy (robtobey2):

(4/3) pi r^3=80

OpenStudy (phi):

I hope you know rob meant (4/3) pi r^3=80 pi

OpenStudy (phi):

to find r^3, divide both sides by pi then multiply both sides by ¾ you should end up with r^3 = some number then take the "cube root" of both sides that will give you r= some number

OpenStudy (phi):

once you know r, you can use the formula for the area of a sphere A= 4 pi r^2

OpenStudy (wolf1728):

Volume = (4/3) * PI * radius^3 80 * PI = (4/3) * PI * radius^3 The 'PI's' cancel out leaving 80 = (4/3) * radius^3 radius^3 = 80/(4/3) radius^3 = 60 radius = cube root (60)

OpenStudy (ellamoyseyuk):

i thought it would be v =4/3 pi(3.93)^3 = v = 4/3 pi(60.7), v = 60.7 x4/3 = 80pi

OpenStudy (ellamoyseyuk):

@triciaal @jim_thompson5910 please help is that the correct way to do it

jimthompson5910 (jim_thompson5910):

r is unknown for now v = 80pi is the given volume

OpenStudy (ellamoyseyuk):

v =4/3 pi(3.93)^3 = v = 4/3 pi(60.7), v = 60.7 x4/3 = 80pi would r =3.93

jimthompson5910 (jim_thompson5910):

\[\Large V = \frac{4}{3}\pi*r^3\] \[\Large 80\pi = \frac{4}{3}\pi*r^3\] \[\Large 80 = \frac{4}{3}*r^3\] \[\Large \frac{3}{4}*80 = \frac{3}{4}*\frac{4}{3}*r^3\] \[\Large 60 = r^3\] \[\Large r^3 = 60\] \[\Large r = \sqrt[3]{60} = 60^{1/3} \approx 3.91486764116887\]

jimthompson5910 (jim_thompson5910):

use \[\Large r \approx 3.91486764116887\] to find the surface area

OpenStudy (ellamoyseyuk):

v = 4/3 pi *r^3 v = 4/3 pi*3.9^3 v = 4/3 pi*59.96 v = 4/3*59.96*pi v =79.95 pi

jimthompson5910 (jim_thompson5910):

you're using the wrong formula

jimthompson5910 (jim_thompson5910):

Surface area of sphere \[\Large SA = 4\pi*r^2\]

OpenStudy (triciaal):

@jim_thompson5910 person did the volume you did the area

OpenStudy (ellamoyseyuk):

S.A = 4p i*3.9^2 S.A = 4 pi*59.319

OpenStudy (ellamoyseyuk):

how would i do the rest

OpenStudy (triciaal):

check your calculation for r^2

jimthompson5910 (jim_thompson5910):

3.9^2 does not turn into 59.319

OpenStudy (ellamoyseyuk):

sorry i put three instead of two

OpenStudy (ellamoyseyuk):

15.21 = 15.21 x 4 = 60.84

jimthompson5910 (jim_thompson5910):

so the approx surface area is 60.84pi = 60.84*3.14 = 191.0376

jimthompson5910 (jim_thompson5910):

rounding that to the nearest whole number gives 191 square cm

OpenStudy (ellamoyseyuk):

okay thank u

jimthompson5910 (jim_thompson5910):

you're welcome